OpenStudy (oliviagggg):

Which of the following solutions would you multiply both sides of the equation by n Solve m/n =p for m Solve m - n = p for m Solve mn = p for m Solve m + n = p for m

OpenStudy (oliviagggg):

I have a feeling that its the second one but I'm not sure

OpenStudy (whpalmer4):

if you multiply by \(n\) you undo division by \(n\)

OpenStudy (whpalmer4):

which of those equations has the target variable (\(m\)) being divided by \(n\)?

OpenStudy (oliviagggg):

The third one?

OpenStudy (whpalmer4):

does does \(mn\) mean divide \(m\) by \(n\)?

OpenStudy (oliviagggg):

None of them have a dividing sign

OpenStudy (whpalmer4):

here, let me show you what you get when you multiply both sides of each equation by \(n\) \[m+n=p\]\[m*n + n*n = p*n\]\[mn + n^2=np\] that didn't get us any closer to having \(m\) alone on one side of the \(=\), did it? that is what we are trying to accomplish when we solve for \(m\)

OpenStudy (whpalmer4):

what does \(m/n\) mean to you?

OpenStudy (oliviagggg):

M*n or m multiplied by n

OpenStudy (whpalmer4):

hint: \[m/n=\frac{m}n=m\div n\]

OpenStudy (oliviagggg):

So is it the first one?

OpenStudy (whpalmer4):

let's try it! \[m/n=p\]\[\frac{m}n=p\]\[n*\frac{m}n=p*n\]\[\cancel{n}*\frac{m}{\cancel{n}}=n*p\]\[m=n*p\] that gives us a solution for \(m\), don't you agree?

OpenStudy (oliviagggg):

Yes so it is the first one because that is all I see that it can be at this point

OpenStudy (oliviagggg):

OK now I get it thank you for the help